OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9 @fateal

OpenStudy (haseeb96):

y=mx +b is the slope intercept form So just multiply 2 with y-3 and

OpenStudy (anonymous):

so the answer is c?

OpenStudy (haseeb96):

2y-6=x+6 2y=x+6+6 2y=x+12 Y=x/2+12/2 y=x/2+6 this is the slope intercept form

OpenStudy (haseeb96):

Yes u r correct.

OpenStudy (anonymous):

thank u!!

OpenStudy (haseeb96):

Most welcome just help the peoplee